### v2

# Seeing through Light and Darkness: Sensor-Physics Grounded Deblurring HDR NeRF from Single-Exposure Images and Events

Abstract
Novel view synthesis from low dynamic range (LDR) blurry images, which are common in the wild, struggles to recover high dynamic range (HDR) and sharp 3D representations in extreme lighting conditions. Although existing methods employ event data to address this issue, they ignore the sensor-physics mismatches between the camera output and physical world radiance, resulting in suboptimal HDR and deblurring results. To cope with this problem, we propose a unified sensor-physics grounded NeRF framework for sharp HDR novel view synthesis from single-exposure blurry LDR images and corresponding events. We employ NeRF to directly represent the actual radiance of the 3D scene in the HDR domain and model raw HDR scene rays hitting the sensor pixels as in the physical world. A pixel-wise RGB mapping field is introduced to align the above rendered pixel values with the sensor-recorded LDR pixel values of the input images. A novel event mapping field is also designed to bridge the physical scene dynamics and actual event sensor output. The two mapping fields are jointly optimized with the NeRF network, leveraging the spatial and temporal dynamic information in events to enhance the sharp HDR 3D representation learning. Experiments on the collected and public datasets demonstrate that our method can achieve state-of-the-art deblurring HDR novel view synthesis results with single-exposure blurry LDR images and corresponding events.
Figure 1: Events corresponding to an LDR blurry image contain spatial difference (High Dynamic Range) and temporal difference (High Temporal Resolution) information of the scene radiance, making it an ideal data for enhancing HDR sharp 3D scene reconstruction. See-NeRF models the physical imaging process with proposed RGB and event mapping fields to bridge the sensor-physics discrepancy, leveraging events to infer scene dynamics and achieving sharper and better HDR novel view synthesis results compared to the physical unaligned method EvHDR-NeRF.
1 Introduction
Recent novel view synthesis works, such as neural radiance fields (NeRF) [31] and 3D gaussian splatting (3DGS) [23], achieve high-fidelity and high-speed novel view synthesis (NVS) with ideal low dynamic range (LDR) sharp images as input, respectively. However, their performance suffers from two limitations: handling extreme lighting scenes with both highlight and low-light areas and robustness to camera motion blur, which are caused by the inherent information loss in the blurred and compressed LDR images with both underexposed and overexposed regions. Some recent image-based NeRF and 3DGS related works utilize multi-exposure LDR images [6, 4, 27] to compensate for the limited dynamic range and model the camera motion [29, 43, 54] to mitigate the influence of motion blur, while they are still susceptible to the requirement of cumbersome multi-exposure capture and severe motion blur in long-exposure conditions.
Event cameras provide a promising sensing modality by asynchronously recording brightness changes in the log domain, thereby preserving high temporal resolution and dynamic range that are typically lost in conventional frame-based cameras [12]. As shown in the upper half of Fig. 1, event data can effectively record spatial-temporal differential information even in blurred, overexposed, and underexposed areas of an LDR image. Motivated by this advantage, recent approaches have explored using event and RGB image (ERGB) data for the deblurring novel view synthesis task [35, 34, 5, 36, 10], while their performance is limited by a lack of grounding in the sensor-physics mismatch between the recorded RGB image and the physical scene brightness. EvHDR-NeRF [6] first incorporates events into the single-exposure HDR and deblurring NeRF reconstruction task. However, its framework largely follows HDR-NeRF [19] by applying the same tone mapping strategy directly to 3D points and extending it to the event-based setting without considering the event generation character.
We argue that achieving accurate and physically consistent 3D representation learning in the ERGB setting requires more than directly extending existing image-based frameworks. In particular, it is essential to explicitly model how scene radiance is transformed into actual sensor measurements through device-specific processes. For RGB cameras, this involves temporal integration and nonlinear camera response functions; for event cameras, this includes contrast thresholding, latency effects, and photometric quantization. Without accounting for these physical transformations, the rendered outputs of the learned model may not align with the supervision signals derived from input images and events, weakening the learning of scene geometry and appearance as shown in the lower half of Fig. 1. Therefore, a sensor-physics grounded modeling of both RGB and event sensing pipelines is critical for fully exploiting the complementary information in ERGB data and achieving robust HDR and deblurring NVS performance.
In this paper, we present a novel framework that exploits Single-Exposure blurry images and corresponding Events to enable sharp HDR NeRF reconstruction (See-NeRF). Specifically, built upon the physical imaging principles of RGB and event sensors, we adopt NeRF to focus on learning the actual radiance of the scene in the HDR domain and introduce two differentiable mapping fields: a pixel-wise RGB mapping field that models exposure integration and camera response, and an event mapping field that accounts for latency and photometric deviations in event generation. These modules align the predicted radiance with the sensor measurements and are jointly optimized with the NeRF network under the supervision of input images and events. By modeling the radiance-to-measurement process explicitly, See-NeRF can accurately reconstruct physically consistent, sharp HDR 3D representations. Experiments on synthetic and real datasets demonstrate that See-NeRF achieves state-of-the-art HDR and deblurring novel view synthesis results. Our contributions are summarized as follows:
1) We propose See-NeRF, a sensor-physics grounded framework that uses events to compensate for the scene dynamics to learn a sharp HDR 3D representation from single-exposure blurry LDR images in extreme lighting scenes.
2) We introduce a 2D pixel-wise RGB mapping field and a latency-aware photometrically calibrated event mapping field to bridge the discrepancy between physical radiance values and sensor measurements.
3) We collect both synthetic and real datasets for training and testing. The results show that our method achieves the best performance on both ERGB-based deblurring and HDR novel view synthesis tasks.
2 Related work
2.1 HDR and deblurring novel view synthesis
Neural radiance fields (NeRF) [31] extensively promote the development of novel view synthesis (NVS) tasks and 3D implicit representation learning. 3D gaussian splatting (3DGS) [23] uses gaussian point clouds to explicitly represent 3D scenes and parallelized rasterization workflow to improve the speed of 3D reconstruction and view rendering. However, NeRF and 3DGS can only learn a sharp LDR scene representation with sharp LDR images. To address the HDR NVS, RawNeRF [30] and RawGS [22] aim to reconstruct an HDR 3D representation from noisy HDR raw images. HDR-NeRF [19] and HDR-GS [4] introduce a 3D points tone mapper to transform the HDR 3D representation into the LDR 3D representation, optimized with multi-exposure LDR images. HDR-HexPlane [46] further extends HDR-NeRF to dynamic scenes. GaussHDR [27] accepts an additional context feature as input to improve HDR NVS. Gaussian in the dark [49] uses a CNN-based feature map to enhance the tone mapper learning. To address the Deblurring NVS from blurry images, Deblur-NeRF [29] proposes a deformable sparse kernel module to model the blur kernels. BAD-NeRF [43] and BAD-GS [54] transform the camera poses into SE(3) space and interpolate the poses to recover the camera motion trajectory. DP-NeRF [24] imposes 3D consistency and refines the color with the relationship between depth and blur. Although these methods have achieved promising performance in HDR and deblurring NVS tasks, they still fundamentally suffer from the inherent information loss of the motion-blurred LDR images.
2.2 ERGB-based novel view synthesis
Event cameras are bio-inspired sensors detecting brightness changes at individual pixels with high temporal resolution [12]. This scene radiance measurement paradigm has higher dynamic range and temporal resolution compared to traditional frame-based RGB cameras, attracting a lot of research in the fields of event-based image deblurring [33, 21, 26, 40, 47], HDR image reconstruction [42, 15, 44, 32, 45, 14, 8], and high-speed HDR video generation [37, 57, 48, 55]. Some recent works explore event-based NVS [20, 39, 28, 28, 29, 11, 53, 13, 51]. While other works integrate both image and event (ERGB) data into the deblurring NVS task, such as NeRF-based works [35, 34, 5, 36] and 3DGS-based works [10, 18, 50] Addressing the limitations of blurry LDR inputs, EvHDR-NeRF [6] first jointly exploits single-exposure images and events to reconstruct high-fidelity HDR radiance fields. However, these methods do not fully consider the gap between sensor output and physical radiance, leading to limited HDR and deblurring NeRF novel view synthesis performance.
3 Background
In this section, we leverage the principles of event and LDR image generation to explain why events can enhance HDR deblurring NVS from single-exposure LDR blurry images.
3.1 From scene radiance to LDR images
In an extreme lighting scene with both highlight and low-light areas, a typical RGB camera sensor records the physical scene radiance through a frame-based integral model, generating a raw image with high dynamic range pixel values. To match human perceptual characteristics and reduce storage costs, the raw image is processed by the image signal processor (ISP), applying non-linear ISO gain, RGB demosaicing, tone mapping, gamma compression, white-balance, and quantization to yield an 8-bit LDR image, which is always modeled by a camera response function (CRF) $f$ in practice [2]:
$\mathcal{I}_{LDR}=f(\int_{t_{start}}^{t_{end}}L(t)dt),$ (1)
where $L(t)$ is the raw scene radiance hitting the sensor at time $t$. $t_{start}$ and $t_{end}$ are the start and end time of exposure. $\mathcal{I}_{LDR}$ is the final color value that we can obtain from the LDR image. This integral model and nonlinear CRF $f$ inherently obscure the temporal dynamics of $L(t)$ and compress the critical HDR spatial radiance differential information in highlight and low-light areas, ineluctably leading to artifacts like motion blur, overexposure (pixel values saturation at 255), and underexposure (pixel values clipping at 0) in the LDR images, respectively.
Figure 2: HDR reconstruction from Events. The first row shows the input data. The second row shows the estimated CRF curves and the dynamic range schematic. For a single-exposure LDR image with exposure time $\Delta t_{2}$, its 8-bit pixel values 0-255 (normalized to 0-1 as $\mathcal{I}_{LDR}$ in the figure) represent a limited range of scene radiance $\ln(L(t))$ (width of the blue box). While traditional methods fuse multi-exposure LDR images with exposure times $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ to compensate for the limited scene brightness representation. We leverage the spatial differential in events to infer the scene radiance of low-light and highlight regions from LDR image ($\Delta t_{2}$), which expands the representation of scene radiance, enabling more accurate CRF estimation curves in Sec. 4.2.
3.2 From scene dynamics to events
Event cameras asynchronously detect the change of the scene radiance hitting the sensor $L(t)$ in the log domain and trigger an event $e(x,y,p,\tau)$ whenever the change reaches the threshold $\Theta$. (x,y) represents the pixel coordinates and $\tau$ is the trigger time. The change direction is encoded in $p\in\{1,-1\}$. This spike model can preserve temporal radiance dynamics and spatial radiance gradient distribution:
$\text{Temporal:}\partial_{t}\log L(x,y,t),$ (2)
$\text{Spatial:}\nabla_{xy}\log L(x,y,t),$
which precisely compensates for the temporal and spatial dynamic loss of blurry compressed LDR images (Sec. 3.1).
3.3 Events for deblurring and HDR NVS
According to Sec. 3.1, the core problem of recovering the sharp HDR scene radiance from LDR blurry images is formulated as the joint estimation of the latent sharp luminance $L(t)$ and the CRF $f$. The temporal radiance dynamics of events in Eq. 2 can be used to estimate the latent sharp $L(t)$ along the time dimension from blurry $\mathcal{I}_{LDR}$ as proved in ERGB-based deblurring works [33, 35]. The spatial dynamics of events in Eq. 2 can provide extra radiance differential information for the low-light and highlight areas and expand the dynamic range of the single-exposure LDR image $\mathcal{I}_{LDR}$, as shown in Fig. 2, with which the network in Sec. 4.2 can obtain a wider range of $\ln(L(t))$ as input and infer a CRF curve that is closer to the ground truth.
Figure 3: The data generation of event and RGB cameras (upper parts) and the pipeline of See-NeRF (lower parts) in an extreme lighting scene. We use NeRF to represent the actual radiance of the scene in the HDR domain. The volume rendering simulates the raw scene radiance rays hitting the sensor physically and obtaining the raw sensor values. The predicted LDR images and events are generated with our proposed RGB and event mapping fields, built upon the data generation model of the sensors. The image loss and event loss are employed to jointly supervise the optimization of the two mapping fields and the NeRF network.
4 Method
In this section, we introduce the See-NeRF framework and theoretically prove how it leverages sensor-physics grounded priors to realize better HDR deblurring and NVS with events and single-exposure blurry LDR images.
4.1 HDR scene representation
Firstly, we use the NeRF network $F_{\theta}$ with parameters $\theta$ to represent the actual radiance $\mathbf{e}$ and density $\sigma$ of a 3D point $\mathbf{o}$ in an extreme lighting scene rather than the camera-processed LDR radiance as in the original NeRF:
$(\textbf{e},\sigma)=F_{\theta}(\gamma_{o}(\textbf{o}),\gamma_{d}(\textbf{d})),$ (3)
where d denotes the observation direction and $\gamma(\cdot)$ is used to transform the input into a higher dimension:
$\gamma_{M}(x)=\{\sin(2^{m}\pi{x}),\cos(2^{m}\pi{x})\}_{m=0}^{M}.$ (4)
Then we use volume rendering to simulate the process of HDR scene rays $\mathbf{o}+\mathbf{d}$ hitting the imaging sensor at pixel (x,y) corresponding to the camera pose $p(t)$ and obtain the latent sharp HDR radiance at time $t$:
$E(x,y,p(t))=\sum_{i=1}^{N}T_{i}(1-\exp(-{\sigma}_{i}{\delta}_{i}))\mathbf{e}_{i},$ (5)
$\text{where}\quad T(i)=\exp(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}).$
$\delta_{i}$ is the distance between adjacent sampled 3D points.
4.2 Physically grounded RGB mapping field
To simulate the integral model of the RGB camera, we discretely sample $b+1$ time points $t_{0},t_{1},...,t_{b}$} within the exposure time $t_{start}$ to $t_{end}$ as in [34], and the generated raw pixel value of the RGB imaging sensor during exposure is:
$\hat{\mathcal{I}}_{HDR}(x,y)=\sum_{i=0}^{b}w_{i}E_{i}(x,y,p(t_{i})),$ (6)
where $w_{i}$ is weight parameters calculated by time points $\{t_{i}\}_{i=1}^{b}$ [34]. To bridge the gap between the raw sensor value and the LDR image calculated by the CRF $f$ in Eq. 1, we used three MLPs $f_{crf}$ to fit this function in the log domain for three color channels separately and obtain the estimated LDR pixels:
$\hat{\mathcal{I}}_{LDR}(x,y)=f_{crf}(\text{ln}(\hat{\mathcal{I}}_{HDR}\Delta{t}_{exp})),$ (7)
where $\Delta{t}_{exp}=t_{end}-t_{start}$ is the exposure time.
Figure 4: Effectiveness of our RGB mapping field.
As shown in Fig. 4, unlike HDR-NeRF applying the CRF function directly on the radiance of 3D points $\mathbf{e}$ in Eq. 3, we employ CRF after volume rendering at the pixel-level for the raw pixel values $\hat{\mathcal{I}}_{HDR}$, aligning with the imaging process in the physical world (Fig. 3), which allows $f_{crf}$ to focus on non-linear color tone mapping learning without being affected by the linear volume rendering, while NeRF $F_{\theta}$ focuses on learning 3D density $\sigma$ and raw radiance $\mathbf{e}$ without being affected by $f_{crf}$. Besides, our blur synthesis is more physically grounded with the HDR pixel values (Eq. 6). Therefore, our proposed RGB mapping field can significantly improve the HDR and deblurring NVS results as demonstrated by the experiment in Sec. 5.4.
4.3 Latency-aware and photometrically calibrated event mapping field
Usually, we can estimate the generated event number during $t_{1}$ and $t_{2}$ at pixel $(x,y)$ with an ideal model:
$B(t_{1},t_{2},x,y)=\text{floor}(\frac{\ln(L(t_{2}))-\ln(L(t_{1}))}{\Theta}),$ (8)
where $\text{floor}(\cdot)$ means taking the integer between the input and zero that is closest to the input. However, the actual event generation may deviate from this ideal model with a significant delay in low-light regions [17], and due to the fixed contrast threshold $\Theta$ in event generation, the minimum detectable brightness change is lower-bounded, leading to quantifiable errors in radiometric estimation. To bridge the gap between actual radiance dynamics and actual event generation, we propose an event mapping field comprising three core components: Bayer pattern adaptation, temporal delay compensation, and photometric quantity calibration.
Bayer pattern adaptation: Like in [5, 6], we input the $b+1$ scene radiance values $\{E(x,y,p(t_{i}))\}_{i=0}^{b}$ with three color channels into an RGGB Bayer pattern to obtain the pixel radiance values of the event sensor $\{E_{ev}(x,y,p(t_{i}))\}_{i=0}^{b}$.
Temporal delay compensation: Since the delay of events is determined by the pixel radiance values of the sensor, we use a compensation function to calculate the coefficient:
$\epsilon_{i}=f_{ev}(E_{ev}(x,y,p(t_{i})).$ (9)
Then, we use a first-order low-pass filter to synthesize the event latency with the estimated coefficient $\epsilon_{i}$:
$L^{i}_{lp}$ $=(1-\epsilon_{i})L^{i-1}_{lp}+\epsilon_{i}E_{ev}(p(t_{i})).$ (10)
We omit (x,y) in Eq. 10 because the above operation is the same for all pixels. We set $L^{0}_{lp}=E_{ev}(t_{0})$ and substitute $\{L^{i}_{lp}\}^{b}_{i=0}$ into the event generation model (Eq. 8) to calculate the estimated number of events:
$\hat{B}^{\prime}(t_{i},t_{i+1})=\text{floor}(\frac{\ln(L^{i+1}_{lp})-\ln(L^{i}_{lp})}{\Theta}).$ (11)
Photometric quantity calibration: Since event acquisition exhibits uncertainty when brightness changes are below the threshold (e.g., minor brightness variations may not trigger events) and our sampled time points $t_{i}$ may not precisely align with the timestamp of the last triggered event at this pixel, we pre-estimate an offset from the input events by using $h(\cdot)$, and then subtract this offset for the final estimated number of events to compensate this discrepancy:
$\hat{B}(t_{i},t_{i+1})=\hat{B}^{\prime}(t_{i},t_{i+1})-h({{B}}(t_{i},t_{i+1})),$ (12)
$h(\cdot)$ is explained in detail in the supplementary materials.
4.4 Loss functions and optimization
We use the LDR image loss and event loss to jointly supervise the optimization of HDR scene radiance representation (Eq. 3) and the two camera mapping fields:
[…]
where $\mathcal{X}$ denotes all pixels of the camera sensor. Then the final loss is defined as:
$\mathcal{L}=\lambda\mathcal{L}_{evs}+\mathcal{L}_{ldr}.$ (15)
We take $\lambda=0.005$ and optimize a coarse and a fine network simultaneously as in [34]. During testing, we can obtain the novel view sharp HDR image with Eq. 6 and the sharp LDR image with novel exposure with Eq. 7. We set $b=4$ as a trade-off between performance and training time as in [35]. Detailed implementation details and evaluation on $b$ are provided in Sec. C and Sec. D of the supplementary.
Table 1: Quantitative results on our synthetic and real dataset for the evaluation of novel view HDR and novel exposure synthesis tasks. The results are the average of all scenes in the dataset. HDR-NeRF+ represents training HDR-NeRF with sharp HDR images obtained by the cascaded ERGB-based image deblurring method EDI [33] and HDR reconstruction method HDRev [48].
Our Synthetic Dataset Our Real Dataset
Input HDR Novel Exposure ($\Delta t_{1}$, $\Delta t_{3}$) HDR Novel Exposure ($\Delta t_{1}$, $\Delta t_{3}$)
Methods Image Event PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Reference HDR-NeRFref Sharp ($\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$) - 26.91 .8906 .1705 29.22 .9247 .1265 26.26 .8716 .1576 33.44 .9618 .0950
RGB-Based HDR NVS HDR-NeRF [19] Sharp ($\Delta t_{2}$) - 17.02 .7384 .3450 24.61 .9178 .1349 10.87 .5473 .4229 22.18 .9013 .1230
HDR-GS [4] Sharp ($\Delta t_{2}$) - 13.54 .6993 .5096 15.13 .7428 .3533 14.18 .6944 .4700 19.23 .8407 .1980
Gaussian-DK [49] Sharp ($\Delta t_{2}$) - 12.74 .6826 .3206 25.87 .9197 .1792 14.33 .7113 .4055 26.41 .8966 .1184
GaussHDR [27] Sharp ($\Delta t_{2}$) - 12.96 .7805 .3015 14.69 .7565 .2154 20.39 .8669 .1723 26.77 .9361 .1771
ERGB-Based Deblurring HDR NVS HDR-NeRF+ Blurry ($\Delta t_{2}$) 16.08 .6436 .3696 17.63 .7178 .2786 18.61 .8092 .2872 15.30 .5936 .3444
EvHDR-NeRF [6] Blurry ($\Delta t_{2}$) 21.73 .8192 .3446 24.37 .8712 .2848 19.00 .6962 .2612 23.76 .8830 .1650
See-NeRF Blurry ($\Delta t_{2}$) 24.13 .8927 .1916 27.57 .9270 .1328 26.49 .8953 .1638 29.11 .9364 .1060
Figure 5: Qualitative results on the synthetic “dogroom” scene and real-world “bear” scene of our datasets for the HDR novel view synthesis task. Our See-NeRF achieves the best results on both under-exposure and over-exposure regions.
Figure 6: Qualitative results on the synthetic “diningroom” (upper) and “bathroom” (lower) scenes for novel exposure synthesis task.
5 Experiments
Since EvHDR-NeRF [6] has not released its datasets, we construct both synthetic and real datasets for the task of HDR novel view synthesis from single-exposure blurry LDR images and corresponding events. We also adopt the public Real-World-Challenge dataset [34] with challenging blurry ERGB input and GT sharp test images to evaluate the deblurring effect of See-NeRF. Due to space limitations, we only briefly describe the collected dataset here and provide further dataset generation and experimental details in the Secs. B, C, and D of the supplementary materials.
5.1 Experiment settings
Synthetic dataset: The synthetic dataset consists of 8 HDR “Blender” scenes from HDR-NeRF [19]. Each scene contains 18 training views and 17 test views. We use Blender [1] and v2e [17] to generate blurry images with exposure time $\Delta t_{2}$ and corresponding events as in [35] for training views. We also provide sharp LDR images with exposure times $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ for the image-based HDR NVS methods without deblurring effects. For test views, we provide GT sharp HDR images and LDR images with $\Delta t_{1},\Delta t_{3}$ for HDR and novel exposure evaluation as in [19].
Real dataset: The real dataset includes 5 scenes captured under extreme lighting conditions using the DAVIS 346 event camera [41]. Each scene has 16 training views and 12 test views. For training views, we use a handheld DAVIS 346 to capture the blurry LDR images with exposure time $\Delta t_{2}$ along with the corresponding event data. We also captured sharp LDR images with a tripod with exposure times $\Delta t_{0},\Delta t_{2},\Delta t_{4}$. For the test views, we use a tripod to capture GT sharp LDR images with exposure times $\Delta t_{1}$, $\Delta t_{3}$. We also capture the GT sharp LDR images with $\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$, and use the classic multi-exposure HDR reconstruction algorithm DeBevec [9] to generate the GT sharp HDR image with the 5 sharp LDR images wtih different exposures for each test view as in RawHDR [56].
Metrics: We use PSNR, SSIM, and LPIPS [52] to quantitatively evaluate the deblurring and novel exposure NVS results. For the HDR NVS task, we tone map [16] the generated GT HDR images into the LDR domain for quantitative and qualitative evaluation as in [4, 19, 56].
5.2 HDR and novel exposure NVS experiment
We choose the image-based HDR NVS works [19, 4, 27, 49] for comparison, and use sharp single-exposure images as input, considering they cannot handle the blurry images. Additionally, we use multi-exposure images as input for HDR-NeRF as a reference. For the ERGB-based methods, we use EDI [33] and HDRev [48] to deblur the blurry LDR images and generate corresponding sharp HDR images with event data, which are input into the HDR-NeRF for training (HDR-NeRF+). We also use the latest ERGB-Based deblurring HDR NeRF methods, EvHDR-NeRF, for comparison.
Quantitative Results: As in Tab. 1, our See-NeRF significantly outperforms the compared methods on both synthetic and real datasets. The HDR reconstruction results on the real dataset are even better than HDR-NeRFref. In comparison, RGB-based HDR NVS methods exhibit constrained performance owing to their dependence on multi-exposure input. Cascaded method HDR-NeRF+ achieves superior results compared to RGB-based methods but underperforms EvHDR-NeRF due to its decoupled pipeline that separately handles 3D reconstruction and HDR sharp image reconstruction, constrained by the inherent bottlenecks of EDI and HDRev. EvHDR-NeRF realizes the second-best overall results, but is significantly worse than our See-NeRF.
Qualitative Results: Fig. 5 displays the qualitative HDR NVS results, and our See-NeRF achieves the best performance. The results of EvHDR-NeRF exhibit slight color deviation and blur artifacts, which stem from inaccurate CRF estimation and unmodeled radiometric integration of the RGB sensor. HDR-NeRF+ generates reconstructions with elevated luminance but significant color distortion, a limitation primarily attributable to HDRev’s suboptimal radiance estimation. Lack of multi-exposure input, RGB-based methods have serious color casts, especially in the overexposed parts of the real data. For the novel exposure synthesis task, See-NeRF perfectly recovers both overexposed and underexposed areas, while other methods are affected by color and brightness mismatch as in Fig. 6.
Table 2: Results on the Real-World-Challenge dataset of the deblurring NVS task. The results are the average of all five scenes.
Methods PSNR↑ SSIM↑ LPIPS↓
RGB-Based Deblurring NVS BAD-NeRF [43] 25.63 .8575 .4400
BAD-GS [54] 27.72 .9039 .3545
DP-NeRF [24] 28.85 .9226 .3015
ERGB-Based Deblurring NVS EBAD-NeRF [36] 27.74 .9010 .3349
Ev-DeblurNeRF [5] 27.83 .9132 .3157
E2GS [10] 28.04 .9211 .2538
E2NeRF [35] 29.92 .9346 .2356
E3NeRF [34] 31.40 .9464 .2000
ERGB-Based Deblurring HDR NVS EvHDR-NeRF [6] 27.19 .8960 .3731
See-NeRF (Ours) 32.70 .9564 .1574
Figure 7: Qualitative results on the “shelf” scene of the Real-World-Challenge data for the deblurring NVS task.
5.3 Deblurring NVS experiment
In this experiment, we compared our method with recent RGB-based and ERGB-based deblurring NVS works, and the quantitative results are shown in Tab. 2. Our See-NeRF achieves significant performance improvements, especially on the LPIPS, proving that our proposed event and RGB mapping fields effectively bridge the difference between the scene radiance and the sensor output, thereby strengthening the use of non-ideal data to reconstruct the sharp 3D scene representation. Fig. 7 shows that See-NeRF successfully recovers the blue eyes and two pink cheek dots of the rabbit, while the best ERGB-based methods E3NeRF miss these subtle details. The results of the best RGB-based method, DP-NeRF, are inferior due to the lack of event enhancement.
Table 3: Quantitative results of our ablation study. The results are the averages of all scenes.
Our Synthesis Data (HDR) Our Real Data (HDR) Real-World-Challenge Data
Event Input Event Mapping Field RGB Mapping Field PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
See-NeRFBase - - - - - - - - - 28.79 .9201 .2939
See-NeRFNoEv - - ours 22.19 .8361 .2600 19.84 .7881 .2502 30.22 .9350 .2425
See-NeRFNoEMF - ours 22.23 .8682 .2254 25.87 .8867 .1701 32.47 .9547 .1686
See-NeRF3D ours 3D Points CRF [19, 4, 6] 20.67 .8561 .2306 24.95 .8323 .1962 31.58 .9534 .1634
See-NeRFEvD eCRF [5] ours 22.02 .8669 .2474 16.43 .6932 .3210 29.52 .9267 .2776
See-NeRF ours ours 24.13 .8927 .1916 26.49 .8953 .1638 32.70 .9564 .1574
5.4 Ablation study
Effect of RGB mapping field: Comparing See-NeRFBase (training the base model without events) and See-NeRFNoe (training See-NeRF without events) in Tab. 3 shows that our RGB mapping field enables the network to perform HDR NVS while improving the deblurring NVS results.
Pixel-Wise CRF versus 3D points CRF: The results of See-NeRF3D of Tab. 3 show that substituting our pixel-wise RGB mapping field with 3D point CRF from [19, 4, 6] significantly reduced performance in HDR and deblurring NVS and leads to incorrect CRF curve estimation in Fig. 8.
Effect of event input: The second and third lines of Tab. 3 prove that event supervision plays a key role in both HDR and deblurring NVS as theoretically deduced in Sec. 3. Fig. 8 further indicates that events can help our model to learn the correct CRF curve of the RGB sensor.
Effect of event mapping field (EMF): Comparing See-NeRFNoEMF and See-NeRF in Tab. 3, our EMF marginally helps align the CRF curve (Fig. 8), achieving further improvement on HDR and Deblurring NVS. However, replacing our EMF with the eCRF model in EvDeblur-NeRF [5] leads to barely satisfactory reconstruction results and CRF estimation (See-NeRFEvD in Tab. 3 and Fig. 8).
Figure 8: Qualitative results (estimated CRF curve and HDR NVS results) on our real “computerhard” scene for ablation study.
Table 4: Discussion on Robustness of See-NeRF.
Robustness Different $\Delta t$ Input Robustness to Event Noise
Single-Exp PSNR↑ SSIM↑ LPIPS↓ Noise Level PSNR↑ SSIM↑ LPIPS↓
$\Delta t_{0}=0.001s$ 24.66 .8886 .2547 $\times 1$ 25.08 .8956 .2324
$\Delta t_{2}=0.016s$ 25.08 .8956 .2324 $\times 2$ 24.99 .8956 .2342
$\Delta t_{4}=0.256s$ 25.68 .9062 .2193 $\times 4$ 24.86 .8948 .2343
5.5 Discussions
Robustness to LDR images with different exposures and event noise: As shown in Tab. 4, we conduct a robustness experiment on the synthetic “Catroom” scene for the HDR NVS task. Firstly, we use the single-exposure images with $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ as inputs, respectively. The results show that single-exposure inputs with longer exposure times yield better results, indicating that the low-light regions of the scene are more difficult to learn than the high-light regions. Secondly, we use events with different noise levels as input. Tab. 4 and the supplement video results on real data show that event noise does not lead to a significant performance degradation, because NeRF training refers to multi-view input, which can inherently weaken the impact of sparse random noise in each individual view, as proved in RawNeRF [30]. Therefore, See-NeRF does not target the processing of event noise, which could be a future optimization direction of ERGB-based NVS methods.
Table 5: Discussion on training time and availability on 3DGS.
Novel Exposure HDR Time
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ hours
HDR-GS 19.19 .7745 .2537 9.25 .4036 .6376 0.03
Ours + 3DGS 20.39 .8204 .2166 18.27 .7244 .3481 0.05
Ours 24.15 .8576 .1154 19.43 .7979 .2469 2.5
Training time and availability on 3DGS: The training time of our See-NeRF (2.5 h) is slightly longer than that of HDR-NeRF (2.1 h) due to the introduction of event-related computation. We also replace the NeRF network of our method with 3DGS and conduct an experiment on the “computerhard” scene of our real data. As in Tab. 5, “Ours + 3DGS” is better than HDR-GS, especially on the HDR NVS task, and the training speed is also increased by about 50 times with the high-speed rendering of 3DGS. However, the performance of “Ours + 3DGS” is significantly lower than See-NeRF on the novel exposure task. We think this is because the MPL-based camera mapping fields do not perfectly match the 3DGS during optimization, and ERGB-based HDR 3DGS may be a promising future work.
6 Conclusion
In this paper, we propose a unified Sensor-Physics grounded framework, See-NeRF, for deblurring HDR NVS from single-exposure LDR images and corresponding events, which bridges the scene radiance and the sensor-recorded ERGB data based on the physical imaging laws and achieves the state-of-the-art deblurring HDR NVS results.
A Details of events calibration and division
Events calibration function $h(\cdot)$: For the input event bin $B(t_{i},t_{i+1})$ between $t_{i}$ and $t_{i+1}$, we use $(x,y,t_{first}^{i},p_{first}^{i})$ and $(x,y,t_{last}^{i},p_{last}^{i})$ to represent the first and last triggered event at pixel $(x,y)$ respectively. Then the offset of event set $B(t_{i},t_{i+1})$ can be expressed as:
$\begin{split}h(B(t_{i},t_{i+1}))=&\phi(\frac{(t_{i}-t_{last}^{i-1})}{t_{first}^{i}-t_{last}^{i-1}}){p_{first}^{i}}-\\ &\phi(\frac{(t_{i+1}-t_{last}^{i})}{t_{first}^{i+1}-t_{last}^{i}}){p_{first}^{i+1}},\end{split}$ (16)
where $\phi(\cdot)$ represents the function that if we can not find the $t_{last}^{i-1}$ or $t_{first}^{i+1}$ (there is no event triggered in the previous event bin or the next event bin at pixel $(x,y)$), the output is $0.5$. Since this operation is the same on all pixels, we omit the $(x,y)$ in Eq. 16.
Events division: For each training view, we split the input event stream $B(t_{start},t_{end})$ that corresponding to the input LDR blurry image $\mathcal{I}_{LDR}$ into $b$ bins $B(t_{i},t_{i+1})\}_{i=0}^{b-1}$, where $t_{0}=t_{start}$ and $t_{b}=t_{end}$ are the start and end exposure time of the image. $t_{1},...,t_{b-1}$ are the time points that divide the event stream into $b$ event bins with an equal number of events as in the E3NeRF [34], which leverages the temporal blur prior in events distribution into event loss optimization to achieve better results.
B Details of our dataset generation
B.1 Synthetic dataset:
Our synthetic dataset consists of 8 HDR Blender [1] scenes (“bathroom”, “catroom”, “dogroom”, “diningroom”, “sofa”, “sponza”, “toyroom”, “warmroom”) in the synthetic data of HDR-NeRF [19], which have both highlight and darkness areas in each scene, resulting in the captured single-exposure images containing both overexposure and underexposure areas as in the first column of Fig. 12. All the data are at a resolution of $400\times 400$. Below, we introduce the training and test data generation process, respectively.
Training images and events generation: For each scene, there are 18 views of LDR blurry images with a single exposure time $\Delta t_{2}$ and corresponding events within the exposure time for training. For each view, we add camera shaking and generate 17 sharp HDR raw images $\{\mathcal{I}_{HDR}^{i}\}_{i=0}^{16}$ at time $t_{i}$ with camera poses $p(t_{i})$. The camera pose $p(t_{i})$ changes over time, indicating the camera shake process. Note that $t_{start}\leq t_{i}\leq t_{end}$, where $t_{start}$ and $t_{end}$ are the start time and end time of the exposure as defined in the main paper. Then the 17 sharp HDR raw images are averaged to obtain the blurred HDR raw image $\mathcal{I}_{HDR}^{blur}$. We used the classic Reinhard tonemapping [38] algorithm in Eq. 17 to convert the blurred HDR raw image into an LDR single-exposure blurred image with exposure time $\Delta t_{2}$ for training, which is the same as in HDR-NeRF [19]:
$\mathcal{I}_{LDR}^{\Delta t_{2}}=(\frac{\Phi\Delta{t_{2}}{\mathcal{I}}_{HDR}^{blur}}{\Phi\Delta{t_{2}}{\mathcal{I}}_{HDR}^{blur}+1})^{\frac{1}{2.2}}.$ (17)
$\Phi$ is a scale factor, and we take it as 62.5. The 17 sharp HDR raw images $\{\mathcal{I}_{HDR}^{i}\}_{i=0}^{16}$ are also input into the event simulator v2e [17] with “noisy” option, which involves the event latency and noise simulation into the data generation. Then the events corresponding to the synthetic LDR blurry image $\mathcal{I}_{LDR}^{\Delta t_{2}}$ is obtained:
$B(t_{start},t_{end})=\mathbf{v2e}(\{\mathcal{I}_{HDR}^{i}\}_{i=0}^{16}).$ (18)
Note that we add a Bayer filter in the v2e to generate the color events as in the DAVIS 346 Color [41] event camera.
Additional training images: For each training view, we also use Eq. 17 to tonemap the 8-th raw sharp image $\mathcal{I}_{HDR}^{7}$ in $\{\mathcal{I}_{HDR}^{i}\}_{i=0}^{16}$ into LDR domain with exposure times $\Delta t_{0}$, $\Delta t_{2}$ and $\Delta t_{4}$, which are used to train the RGB-based HDR NVS works [19, 4, 27, 49] without deblurring effects and to train HDR-NeRF [19] with multi-exposure sharp LDR images as input for reference.
Test images generation: For each scene, there are 17 novel views for testing. For each test view, we render 1 ground-truth (GT) sharp HDR raw image $\mathcal{I}_{HDR}^{sharp}$ with Blender for the HDR novel view synthesis (NVS) test. We also use Eq. 17 to transform the sharp HDR image $\mathcal{I}_{HDR}^{sharp}$ into two GT sharp LDR images at novel exposure times $\Delta t_{1}$ and $\Delta t_{3}$ for the novel exposure NVS test.
B.2 Real dataset:
Our synthetic dataset consists of 5 real-world extreme lighting scenes (“bear”, “computer”, “computerhard”, “desktop”, “table”) with both highlight and darkness areas, resulting in the captured single-exposure LDR images containing both overexposure and underexposure areas. The upper part of Tab. 6 shows the minimum and maximum luminance of the scenes in our real data, which contains a very wide dynamic range. All data is obtained by a DAVIS 346 Color camera [41], which can capture spatial-temporal aligned images and events at resolution $346\times 260$.
Training images and events generation: For each scene, there are 16 training views. For each view, we capture 1 LDR blurry image at a single exposure time $\Delta t_{2}$ and corresponding events within the exposure time with a handheld DAVIS 346 Color camera.
Additional training images: For each training view, we also capture three sharp LDR images at exposure times $\Delta t_{0}$, $\Delta t_{2}$, and $\Delta t_{4}$ with a tripod-fixed DAIVS 346 Color camera to train the RGB-based HDR NVS works [19, 4, 27, 49] without deblurring effects and to train HDR-NeRF [19] with multi-exposure sharp LDR images as input for reference.
Test images generation: For each scene, there are 12 novel views for testing. For each test view, we capture 5 LDR sharp images at exposure times $\{\Delta t_{i}\}_{i=0}^{4}$ with a tripod-fixed DAIVS 346 Color camera. The LDR sharp images with novel exposure times $\Delta t_{1}$ and $\Delta t_{3}$ are used for the novel exposure NVS test, which is the same as the synthetic dataset. The GT HDR sharp image for the HDR NVS test is obtained by merging the five LDR sharp images with exposure times $\{\Delta t_{i}\}_{i=0}^{4}$ with the classic multi-exposure HDR reconstruction algorithm Debevec [9] as in RawHDR[56].
Figure 9: Additional evaluation of the HDR NVS performance boundaries between multi-exposure image-based HDR-NeRF and our ERGB-based See-NeRF on the real “Table” scene. The upper and lower parts of the figure are the estimated CRF curve and HDR NVS results, respectively. See-NeRF leverages events to replace the CRF curve correction effect of the LDR images with exposure times $\Delta t_{0}$ and $\Delta t_{4}$ as in HDR-NeRFref. Besides, See-NeRF is robust to the image motion blur compared to HDR-NeRFref-blur.
Table 6: The exposure time settings of our synthetic and real datasets and the illumination conditions of our real-world scenes.
Synthetic Real-World Scenes
Scenes Bear Computer Computerhard Desktop Table
Min Scene luminance - 70 lux 50 lux 30 lux 20 lux 10 lux
Max Scene luminance - 2000 lux 4000 lux 5000 lux 1000 lux 1500 lux
Exposure Time $\Delta{t}_{0}$ 0.001 s 0.05 s 0.02 s 0.002 s 0.05 s 0.05 s
Exposure Time $\Delta{t}_{1}$ 0.004 s 0.10 s 0.04 s 0.004 s 0.10 s 0.10 s
Exposure Time $\Delta{t}_{2}$ 0.016 s 0.15 s 0.06 s 0.006 s 0.15 s 0.15 s
Exposure Time $\Delta{t}_{3}$ 0.064 s 0.20 s 0.08 s 0.008 s 0.20 s 0.20 s
Exposure Time $\Delta{t}_{4}$ 0.256 s 0.25 s 0.10 s 0.010 s 0.25 s 0.25 s
C Implementation details
Our framework is built upon E2NeRF [35], and all experiments are conducted on a single RTX 3090 GPU. We set the number of sampled points on each ray to 64 and 128 for the coarse and fine networks, respectively. The batch size of sampled rays is set to 512 and 1024 for the real data and synthetic data, respectively, and we take 50k iterations for each scene. The above NeRF network settings are the same for all the compared NeRF-based methods and our See-NeRF in the experiments. For the 3DGS-based methods, the number of iterations is set to 30000, and other parameters are set to their default parameters for each method. For the detailed evaluation of hyperparameters $b$, $\lambda$, and training time of our See-NeRF, please refer to Sec. D.2.
C.1 Exposure time settings
The lower part of Tab. 6 shows the value of exposure times $\{\Delta t_{i}\}_{i=0}^{4}$ of our synthetic and real datasets. For the Real-World-Challenge dataset [34] of the deblurring experiment, the exposure times of the input blurry images are 0.1 s, 0.12 s, 0.1 s, 0.1 s, 0.15 s, for the “corridor”, “lab”, “lobby”, “shelf”, “table” scenes, respectively. We use these exposure times to train and test our See-NeRF to obtain the corresponding deblurred novel view LDR images.
C.2 Pose settings
We use the pose generation framework of E2NeRF [35] to obtain the poses $\{p(t_{i})\}_{i=0}^{b}$ corresponding to the blurry images for all the compared methods and our See-NeRF, which ensures the input poses are the same, so the reconstruction result is only affected by the network architecture.
C.3 Evaluation of HDR results
For the HDR NVS task, we first use the “Enhancer” and “Compressor” options of Photomatrix Pro [16] for our synthetic and real data, respectively, to tonemap the reconstructed and ground truth HDR images into the LDR domain. Then we use PSNR, SSIM, and LPIPS [52] to quantitatively and qualitatively show the results in our main paper and supplementary materials, which are the same as [4, 19].
D Additional evaluations
D.1 Single-exposure vs. multi-exposure
In this experiment, we train HDR-NeRF [19] with different input exposure ranges to evaluate the HDR NVS performance boundaries of multi-exposure image-based HDR-NeRF and our single-exposure ERGB-based See-NeRF.
Table 7: Additional evaluation of the HDR NVS performance boundaries between multi-exposure image-based HDR-NeRF and our ERGB-based See-NeRF on the real “Table” scene.
Real “Table” Scene of Our Dataset
Input Data HDR Results
Exposure Time Image Event PSNR↑ SSIM↑ LPIPS↓
HDR-NeRFref $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ Sharp - 28.23 .9207 .1550
[…]
HDR-NeRFref-blur $\Delta t_{0},\Delta t_{2},\Delta t_{4}$ Blurry - 25.54 .8640 .3017
See-NeRF $\Delta t_{2}$ Blurry 27.12 .9093 .1548
Quantitative results: HDR-NeRFref in Tab. 7 demonstrates that multi-exposure image-based HDR-NeRF can only fully reconstruct sharp HDR scenes when the input LDR images are sharp and cover exposure times at all three scales: low ($\Delta t_{0}$), medium ($\Delta t_{2}$), and high ($\Delta t_{4}$). When the input long-exposure images are motion-blurred, the results of HDR-NeRFref-blur will be significantly affected by the motion blur and degraded. In contrast, our See-NeRF achieves the best HDR NVS results with blurry single-exposure LDR images and corresponding events
[…]
Therefore, See-NeRF effectively leverages the spatial and temporal differential information in the low-light, highlight, and blurry regions that are contained in the events to infer the actual brightness of the scene and eliminates the effects of the blurred input, respectively, achieving the best deblurring HDR NVS results with single-exposure input and events, as deduced in the Sec. 3 of the main paper.
Table 8: Evaluation of parameter $b$ and training time.
HDR (“Desktop” scene) Deblurring (“Lab” scene) Time
$b$ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ hours
2 27.10 .9127 .1930 33.98 .9657 .1458 1.75
4 28.13 .9152 .1796 34.56 .9681 .1374 2.50
6 26.89 .8856 .1842 35.21 .9712 .1277 3.23
8 26.49 .8753 .1834 35.57 .9742 .1193 3.95
D.2 Parameters and training time evaluation
Parameter $b$: $b$ in Eq.(6) represents the discretization of Eq.(1) with $b$ discrete time points. Therefore, a larger $b$ will lead to better simulation of image formation and generate better results, but at the same time require more iterations and increase the training time as shown in Tab. 8. For the HDR NVS task, when $b=4$, See-NeRF gets the best results. For the deblurring NVS task, when $b>6$, the improvement is marginal. So we set $b=4$ for the HDR and novel exposure NVS experiments on our proposed datasets and $b=6$ for the deblurring NVS experiments for the Real-World-Challenge dataset.
Training time: Compared to HDR-NeRF (2.1h), See-NeRF (2.5h) increases computation and training time by 19.05% under the same conditions (same NeRF-related parameters, number of iterations, and batch size; on a single RTX 3090 GPU; on our real dataset). The increased computation is mainly due to the introduction of events and calculations related to event mapping fields. However, since no events are involved during test time, the computation and rendering speed are close to HDR-NeRF (At 346*260 resolution, rendering each frame takes 2.5 seconds).
Figure 10: Evaluation of parameter $\lambda$.
Parameter $\lambda$: As shown in Fig. 10, we evaluate the impact of $\lambda$ in real-world “desktop” and “lab” scenes for the HDR and deblurring NVS tasks, respectively. When $\lambda=0.005$, See-NeRF can achieve the best overall results on both tasks. Note that $\lambda=0$ represents train See-NeRF without event loss supervision.
Figure 11: Analysis of sensor-physics grounded deblurring effect.
D.3 Sensor-Physics grounded deblurring effect
Our See-NeRF is significantly better than previous ERGB-based deblurring NVS methods E2NeRF [35], E3NeRF [34], E2GS [10], and Ev-DeblurNeRF [5] as in Tab. 16. This is because these methods all perform the motion blur and events synthesis directly on the nonlinear LDR images rendered by their NeRF network, as shown in Fig. 11. In contrast, our See-NeRF performs the motion blur and events synthesis first on the linear raw HDR values of the scene radiance rendered by the NeRF network, and then tonemaps the blurry raw HDR image to the LDR domain, aligning with the physical imaging process of sensors in the real world. On the other hand, our event mapping field eliminates the effects of event delay and further improves the performance. Therefore, See-NeRF achieves SOTA results in the ERGB-based deblurring NeRF task.
Table 9: Discussion with EvHDR-GS
Imaging Model Event Real Data HDR Real Data Deblurring
Method Integration CRF Non-Idealities PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
EvHDR-GS N/A $\mathcal{F}(x)=x^{a}$ N/A 18.94 .7124 .2510 28.69 .9174 .2674
See-NeRF Eq. (6) MLP-Based RGB CRF Sec. 4.3 26.49 .8953 .1638 32.70 .9564 .1574
See-NeRF# N/A $\mathcal{F}(x)=x^{a}$ Sec. 4.3 19.75 .7239 .2696 28.35 .9242 .2657
D.4 Discussion with EvHDR-GS
A recent work, EvHDR-GS [7], uses 3D gaussian splatting as a backbone for the ERGB-based HDR video reconstruction task. Though it uses a similar design that RGB CRF is applied after volumetric rendering, it does not consider the image integration process and simply models the RGB CRF with one exponential function $\mathcal{F}(x)=x^{a}$ during the 2D mapping, while See-NeRF address this two aspects with Eq. (6) in the main paper and a three-independent MLP-Based RGB CRF model as shown in the Fig. 3 of the main paper. Besides, we model the event non-idealities with our event mapping fields. Therefore, EvHDR-GS underperforms See-NeRF on both deblurring and HDR tasks, and applying its imaging model to See-NeRF also degrades the performance significantly, as shown in Tab. 9.
D.5 Pose estimation
We use the event-enhanced pose estimation strategy of E2NeRF for our See-NeRF. Recent image-based deblurring NVS works provide different camera pose estimation strategies [43, 54, 24] and pose change estimation strategies for dynamic scenes [3, 25]. Integrating the event supervision into these advanced pose estimation strategies, as in EBAD-NeRF [36], would be a promising future research direction for the ERGB-based HDR deblurring NVS task.
E Detailed results
E.1 Detailed quantitative results
Our synthetic dataset: Tab. 10 and Tab. 11 illustrate the detailed quantitative novel exposure NVS results on each scene of our synthetic dataset. See-NeRF achieves the best overall performance. In certain scenes, the novel exposure results of the compared RGB-based HDR NVS methods are better than those of See-NeRF because it uses sharp images as input. Besides, the detailed quantitative results in Tab. 12 and Tab. 13 illustrate that their performance drops significantly on the HDR NVS task, indicating that they are overfit to the novel exposure task, thus learning a wrong scene HDR representation. In contrast, See-NeRF achieves the best HDR NVS results for all three metrics.
Our real dataset: Tab. 14 and Tab. 15 illustrate the detailed quantitative novel exposure and HDR NVS results on each scene of our real dataset. Some results of the compared RGB-based HDR NVS methods are better than those of See-NeRF on LPIPS due to the sharp images input, while our See-NeRF still achieves the best overall performance. For the HDR experiment, See-NeRF is even better than HDR-NeRFref in “Bear”, “Computerhard”, and “Desktop” scenes, resulting in the average results in Tab. 1 of the main paper outperforms HDR-NeRFref. This is likely attributable to the low imaging quality of the DAVIS 346 RGB sensor, which is not suitable for image-based methods.
Real-World-Challenge dataset: Tab. 16 illustrates the detailed quantitative deblurring NVS results on each scene of the Real-World-Challenge dataset [34]. Our See-NeRF realizes the SOTA performance compared to the previous RGB-based and ERGB-based deblurring NVS methods.
E.2 Detailed qualitative results
Our synthetic dataset: Fig. 12 shows the detailed qualitative HDR NVS results on our synthetic dataset, and the results of See-NeRF are closest to GT across all scenes.
Our real dataset Fig. 13 shows the detailed qualitative HDR NVS results on our real dataset, and the results of See-NeRF are closest to GT across all scenes.
Real-World-Challenge dataset Fig. 14 shows the detailed qualitative deblurring NVS results on the Real-World-Challenge dataset. See-NeRF shows impressive deblurring performance compared to other methods.
E.3 Supplementary video results
We also provide supplementary video rendering results to show the impressive HDR and Deblurring NeRF novel view synthesis results of See-NeRF and robustness to noise events of our framework. We highly recommend that the reviewers refer to the attached video for details.
Table 10: Detailed quantitative novel exposure NVS results on each scene of our synthetic dataset.
Input Bathroom Catroom Diningroom Dogroom
Methods Image Event PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Reference HDR-NeRFref Sharp ($\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$) - 29.69 .9148 .1798 26.56 .9089 .1589 29.02 .9493 .1053 25.20 .8537 .1028
RGB-Based HDR NVS HDR-NeRF [19] Sharp ($\Delta t_{2}$) - 27.13 .9161 .1920 24.56 .8954 .1769 23.68 .9454 .1002 21.11 .8382 .1088
HDR-GS [4] Sharp ($\Delta t_{2}$) - 17.17 .7536 .3440 15.66 .7041 .3088 12.73 .6970 .4790 13.24 .7189 .3377
Gaussian-DK [49] Sharp ($\Delta t_{2}$) - 29.99 .9599 .1580 28.46 .9478 .1636 21.21 .9142 .1409 24.24 .9714 .1593
GaussHDR [27] Sharp ($\Delta t_{2}$) - 15.30 .7288 .1925 14.37 .7367 .1858 12.73 .7483 .2230 12.03 .7115 .2473
ERGB-Based Deblurring HDR NVS HDR-NeRF+ Blurry ($\Delta t_{2}$) 19.50 .7311 .3197 18.58 .7288 .2970 14.59 .7336 .2699 14.74 .5893 .3586
EvHDR-NeRF [6] Blurry ($\Delta t_{2}$) 26.43 .8729 .3124 25.11 .8855 .3249 20.45 .8412 .2847 22.81 .8229 .2903
See-NeRF Blurry ($\Delta t_{2}$) 28.13 .9141 .1596 27.45 .9214 .1774 23.28 .9315 .1314 24.81 .8678 .1430
Table 11: Detailed quantitative novel exposure NVS results on each scene of our synthetic dataset.
Input Sofa Sponza Toyroom Warmroom
Methods Image Event PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Reference HDR-NeRFref Sharp ($\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$) - 30.66 .9292 .1139 31.49 .9555 .1161 31.69 .9641 .0694 29.47 .9224 .1657
RGB-Based HDR NVS HDR-NeRF [19] Sharp ($\Delta t_{2}$) - 20.48 .9199 .1248 29.48 .9504 .1252 25.48 .9569 .0934 24.97 .9200 .1583
HDR-GS [4] Sharp ($\Delta t_{2}$) - 12.97 .7777 .2391 18.78 .7352 .4568 15.04 .7682 .3198 15.42 .7873 .3411
Gaussian-DK [49] Sharp ($\Delta t_{2}$) - 24.46 .9787 .1351 20.27 .6563 .3864 28.63 .9681 .1526 29.74 .9609 .1378
GaussHDR [27] Sharp ($\Delta t_{2}$) - 13.79 .7391 .2252 22.34 .8678 .2037 12.49 .7607 .2293 14.50 .7589 .2164
ERGB-Based Deblurring HDR NVS HDR-NeRF+ Blurry ($\Delta t_{2}$) 18.22 .7045 .2436 18.06 .6032 .2696 17.18 .8291 .1948 20.15 .8232 .2753
EvHDR-NeRF [6] Blurry ($\Delta t_{2}$) 24.04 .9048 .2283 24.40 .8551 .2655 25.70 .9142 .2871 26.05 .8728 .2850
See-NeRF Blurry ($\Delta t_{2}$) 28.03 .9399 .1075 33.33 .9693 .0929 28.76 .9533 .0937 26.75 .9185 .1572
Table 12: Detailed quantitative HDR NVS results on each scene of our synthetic dataset.
Input Bathroom Catroom Diningroom Dogroom
Methods Image Event PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Reference HDR-NeRFref Sharp ($\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$) - 25.96 .8412 .2413 26.09 .8914 .2015 25.45 .9280 .1432 24.65 .8332 .1261
RGB-Based HDR NVS HDR-NeRF [19] Sharp ($\Delta t_{2}$) - 19.06 .7886 .3384 19.84 .8531 .2722 11.15 .4079 .6993 14.65 .6379 .3291
HDR-GS [4] Sharp ($\Delta t_{2}$) - 12.16 .6562 .5850 13.58 .7421 .5033 12.90 .6761 .4857 14.56 .6995 .4755
Gaussian-DK [49] Sharp ($\Delta t_{2}$) - 13.57 .7088 .2689 12.04 .7207 .2380 10.71 .5619 .4420 12.87 .7302 .2583
GaussHDR [27] Sharp ($\Delta t_{2}$) - 13.39 .7976 .1398 10.81 .7321 .1784 12.27 .7389 .1540 10.86 .7789 .1763
ERGB-Based Deblurring HDR NVS HDR-NeRF+ Blurry ($\Delta t_{2}$) 16.18 .6439 .3847 14.29 .6153 .3434 12.95 .5215 .6971 16.15 .5081 .3807
EvHDR-NeRF [6] Blurry ($\Delta t_{2}$) 22.85 .7915 .3624 23.56 .8594 .3784 23.11 .8841 .2452 17.62 .7442 .3509
See-NeRF Blurry ($\Delta t_{2}$) 20.83 .8282 .2801 25.08 .8956 .2324 24.71 .9219 .1414 21.99 .8638 .1757
Table 13: Detailed quantitative HDR NVS results on each scene of our synthetic dataset.
Input Sofa Sponza Toyroom Warmroom
Methods Image Event PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Reference HDR-NeRFref Sharp ($\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$) - 28.34 .9000 .1504 26.24 .8841 .2189 30.66 .9525 .0994 27.94 .8948 .1828
RGB-Based HDR NVS HDR-NeRF [19] Sharp ($\Delta t_{2}$) - 13.45 .7020 .3381 24.93 .8736 .2469 14.60 .8315 .2605 18.50 .8126 .2752
HDR-GS [4] Sharp ($\Delta t_{2}$) - 12.29 .6713 .5124 17.00 .7045 .5966 13.30 .7606 .4901 12.53 .6837 .4284
Gaussian-DK [49] Sharp ($\Delta t_{2}$) - 15.00 .7883 .2829 11.20 .5116 .5422 12.85 .7376 .2544 13.66 .7014 .2782
GaussHDR [27] Sharp ($\Delta t_{2}$) - 13.89 .8521 .1467 16.48 .7860 .2726 12.92 .8112 .1398 13.03 .7470 .1647
ERGB-Based Deblurring HDR NVS HDR-NeRF+ Blurry ($\Delta t_{2}$) 17.93 .5986 .3015 17.90 .8002 .2967 17.77 .7607 .2310 15.45 .7002 .3215
EvHDR-NeRF [6] Blurry ($\Delta t_{2}$) 20.63 .8312 .2983 20.91 .7667 .4215 21.24 .8450 .3948 23.93 .8315 .3056
See-NeRF Blurry ($\Delta t_{2}$) 23.55 .9117 .1447 26.74 .9102 .2127 26.01 .9357 .1604 24.13 .8746 .1850
Table 14: Detailed quantitative novel exposure NVS results on each scene of our real dataset.
Input Bear Computer Computerhard Desktop Table
Methods Image Event PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
HDR-NeRFref Sharp ($\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$) - 32.89 .9611 .1208 28.95 .9451 .0675 32.69 .9685 .0440 35.71 .9612 .1353 36.97 .9732 .1072
HDR-NeRF Sharp ($\Delta t_{2}$) - 20.05 .9138 .1498 20.95 .8721 .1098 20.44 .8787 .0977 24.65 .9102 .1381 24.82 .9315 .1196
HDR-GS Sharp ($\Delta t_{2}$) - 18.35 .8736 .1901 18.76 .7913 .1863 19.19 .8257 .1537 21.03 .8965 .2218 18.83 .8163 .2382
Gaussian-DK Sharp ($\Delta t_{2}$) - 25.32 .9541 .1201 23.50 .8821 .1085 21.83 .7411 .1257 30.79 .9440 .1348 30.61 .9619 .1030
GaussHDR Sharp ($\Delta t_{2}$) - 25.07 .9454 .1862 24.63 .9236 .1777 23.28 .9101 .1763 30.13 .9454 .1829 30.74 .9561 .1626
HDR-NeRF+ Blurry ($\Delta t_{2}$) 17.13 .7716 .3123 13.30 .5075 .2917 12.60 .3793 .4650 16.93 .6431 .3710 16.55 .6665 .2819
EvHDR-NeRF Blurry ($\Delta t_{2}$) 22.14 .9059 .1937 23.91 .8760 .1228 21.12 .8187 .1598 25.94 .8984 .1777 25.68 .9160 .1709
See-NeRF Blurry ($\Delta t_{2}$) 30.24 .9591 .1202 27.04 .9481 .0713 24.15 .8576 .1154 31.84 .9536 .1155 32.27 .9621 .1076
Table 15: Detailed quantitative HDR NVS results on each scene of our real dataset.
Input Bear Computer Computerhard Desktop Table
Methods Image Event PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
HDR-NeRFref Sharp ($\Delta t_{0}$, $\Delta t_{2}$, $\Delta t_{4}$) - 30.79 .9479 .1045 28.08 .8976 .1226 17.03 .6914 .2024 27.15 .9006 .2036 28.23 .9207 .1550
HDR-NeRF Sharp ($\Delta t_{2}$) - 10.15 .4789 .4333 10.28 .4684 .4552 9.76 .4989 .4850 11.71 .6101 .3860 12.47 .6804 .3551
HDR-GS Sharp ($\Delta t_{2}$) - 17.11 .8773 .3282 14.29 .5777 .5406 9.25 .4036 .6376 15.00 .7993 .4103 15.24 .8140 .4334
Gaussian-DK Sharp ($\Delta t_{2}$) - 14.27 .8181 .3253 12.75 .6314 .4308 11.09 .4588 .5935 16.47 .8113 .3343 17.08 .8367 .3436
GaussHDR Sharp ($\Delta t_{2}$) - 19.55 .9288 .1939 16.35 .8305 .2451 17.40 .7771 .1689 24.21 .8934 .1432 24.43 .9047 .1102
HDR-NeRF+ Blurry ($\Delta t_{2}$) 21.10 .8996 .2196 18.09 .7855 .2772 12.18 .6128 .4355 18.68 .8443 .3110 23.01 .9039 .1925
EvHDR-NeRF Blurry ($\Delta t_{2}$) 18.17 .7972 .2138 23.18 .8251 .2216 10.39 .1939 .4011 21.85 .8000 .2465 21.40 .8649 .2229
See-NeRF Blurry ($\Delta t_{2}$) 31.91 .9558 .0999 25.85 .8982 .1379 19.43 .7979 .2469 28.13 .9152 .1796 27.13 .9093 .1548
Table 16: Detailed quantitative deblurring NVS results on each scene of the Real-World-Challenge dataset.
Corridor Lab Lobby Shelf Table
Methods PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
RGB-Based Deblurring NVS BAD-NeRF 26.24 .8950 .4761 28.86 .9125 .2386 23.06 .8175 .6265 29.61 .8820 .2487 20.39 .7804 .6102
DP-NeRF 29.64 .9508 .3361 31.40 .9452 .2350 27.52 .9168 .3093 29.92 .9036 .2937 25.79 .8964 .3336
ERGB-Based Deblurring NVS EBAD-NeRF 28.74 .9288 .4018 30.10 .9354 .2052 26.66 .8867 .4048 29.88 .8965 .2646 23.34 .8574 .3981
Ev-DeblurNeRF 31.27 .9546 .2960 31.64 .9570 .1790 22.48 .8611 .4917 31.53 .9285 .2141 22.23 .8647 .3975
E2GS 23.83 .8699 .5516 32.84 .9571 .2131 27.92 .9132 .4155 31.81 .9205 .2348 26.91 .9218 .2807
E2NeRF 30.13 .9584 .2228 32.16 .9581 .1717 28.94 .9258 .2635 31.23 .9109 .2578 27.15 .9196 .2622
E3NeRF 31.50 .9636 .2213 34.02 .9641 .1505 30.25 .9339 .2479 32.48 .9342 .1736 28.75 .9362 .2067
ERGB-Based Deblurring HDR NVS EvHDR-NeRF 29.98 .9444 .3701 27.34 .9078 .3400 27.24 .9064 .3544 28.08 .8639 .3345 23.32 .8575 .4665
See-NeRF 34.12 .9749 .1313 35.21 .9712 .1277 31.12 .9449 .2248 34.40 .9545 .1065 28.64 .9367 .1965
Figure 12: Detailed qualitative HDR NVS results on each scene of our synthetic dataset.
Figure 13: Detailed qualitative HDR NVS results on each scene of our real dataset.
Figure 14: Detailed qualitative deblurring NVS results on each scene of the Real-World-Challenge dataset.